# Near-threshold photon-proton production of $J/\psi$ and $\Upsilon$

### v1

@@ III Gluon GPDS in the Spectator Model
In this section, we present an analytical calculation of the gluon GPDs $H^{g}(x,\xi,t)$ and $E^{g}(x,\xi,t)$ at nonzero skewness $\xi$ in the proton using a spectator model [68]. A variant of this model, incorporating a spectral function for the spectator mass, has previously been used to compute twist‑2 gluon TMD distributions [69, 70]. For the following discussion, it is convenient to introduce the polarization states of the proton in an arbitrary direction:
@@
The gluon distributions in $x$ satisfy the Bose symmetry [66], leading to $H(x,\xi,t)=-H(-x,\xi,t)$ and $E(x,\xi,t)=-E(-x,\xi,t)$. As shown in Eq. (10), the integrals of gluon GPDs over $x$ are restricted to $x\geq 0$, only the two poles at $k_{2}^{-}$ and $k_{3}^{-}$ in Eqs. (28-29) need to be considered,
@@
After selecting the appropriate helicity combinations and carrying out the $k^{-}$ integrals in Eq. (26), the twist‑2 gluon GPDs can be constructed as
@@ IV.1 Parameter determination and derived quantities
The present spectator model adopted in the calculation is similar to that in Ref. [69] except we refrain to include a spectral function for the spectator mass. The model contains four free parameters: the normalization constants $\kappa_{1,2}$; the cut‑off $\Lambda_{X}$ which regulates the large $\bm{k}_{\perp}$-behavior; and the spectator mass $M_{X}$ which is taken larger than the proton mass ($M_{X}>M$). To describe the $x$-dependence of the distributions reliably, we determine these parameters by simultaneously fitting the unpolarized PDF (Eq. 39) and the helicity PDF (Eq. 40) to the next‑to‑leading‑order (NLO) NNPDF3.1 parametrization of $f_{1}^{g}(x)$ [80] and the NLO NNPDFpol1.1 parametrization of $g_{1L}^{g}(x)$ [81] at the scale $\mu_{0}=2\;\text{GeV}$. The fitted results and their uncertainties are listed in Table 1.
@@
Figure 6: Differential cross sections for $\gamma p\rightarrow J/\psi\,p$ as functions of t-t at different average photon energies $8.2\;\text{GeV}\leq E_{\gamma}\leq 11.8\;\text{GeV}$ compared with data from GlueX19 [10] (triangles), GlueX23 [11] (circles), and $J/\psi$‑007 [12] (squares) . The bands depict the parameter uncertainties.
@@
We first analyze all available data for near-threshold $J/\psi$ photoproduction from the GlueX [10, 11] and $J/\psi$‑007 experiments [12]. These data include differential cross sections in 15 photon-energy bins and total cross sections, with the GlueX23 measurements on differential cross section covering the full near-threshold kinematic region. They have been analyzed in recent theoretical studies on gluonic properties of the proton [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. For the parameters related to the $J/\psi$ production [18], we adopt $\alpha_{\text{EM}}=1/137$, $\alpha_{S}=0.3$, $e_{c}=2/3$, $M=0.938\;\text{GeV}$, $M_{J/\psi}=3.097\;\text{GeV}$, and the $J/\psi$ wave function at the origin [84, 85]
@@
In Fig. 6 we plot the model results (solid curves) for the differential cross sections as functions of t-t at different $\langle E_{\gamma}\rangle$ compared with the experimental data measured at different t-t bins. The triangle, circle, and square data points corresponds to the data from GlueX19 [10] , GlueX23 [11] , and $J/\psi$‑007 [12] , respectively. The blue bands depict the uncertainties from the model parameters. Overall, the model describes the data reasonably well, with slight deviations at $\langle E_{\gamma}\rangle=9.86$ and $10.82\;\text{GeV}$. The differential cross sections exhibit the characteristic diffractive peak at forward $t$ and an exponential fall-off at larger momentum transfer. The GlueX23 data (circles) span the entire physical $t$ range; at the lowest average energy $\langle E_{\gamma}\rangle=8.93\;\text{GeV}$ an enhancement appears at the largest $|t|$, possibly signaling contributions beyond the pure $t$-channel mechanism (e.g., $u$- or $s$-channel exchanges) [23, 86]. While our model does not reproduce this enhancement at $8.93\;\text{GeV}$, similar features emerge at $\langle E_{\gamma}\rangle=9.86$ and $10.82\;\text{GeV}$.
@@
In Fig. 8 we show the predicted differential cross sections near threshold ($\langle E_{\gamma}\rangle=57.2\;\text{GeV}$, left) and farther above threshold ($\langle E_{\gamma}\rangle=70\;\text{GeV}$, right). The total cross section is displayed in Fig. 9. The larger bottom-quark mass makes the heavy-quark limit more reliable for $\Upsilon$ production, reducing theoretical uncertainties from higher=order corrections. However, the production rate is strongly suppressed by the heavy quark mass, leading to cross sections much smaller than that for $J/\psi$. Verification of these results awaits more precise future measurements.
@@ V Summary
In this work we studied the near-threshold exclusive photoproduction of heavy quarkonia within the GPD formalism, employing a spectator model for the proton. The amplitudes of these processes are dominated by the leading moments of gluon GPDs, offering a probe of gluon distributions at large skewness $\xi$. The model parameters were determined by simultaneously fitting $f_{1}^{g}(x)$ from the NNPDF3.1 parameterization and F $g_{1L}^{g}(x)$ from the NNPDFpol1.1 parameterization at $\mu_{0}=2\;\text{GeV}$. The gluon GPDs $H^{g}(x,\xi,t)$ and $E^{g}(x,\xi,t)$ at nonzero $\xi$ were calculated analytically at tree level via two-gluon exchange, separately in the regions $0\leq x\leq\xi$ and $\xi<x\leq 1$.
Several gluon-related quantities were extracted and compared with recent lattice results. The gluonic contribution to the quantum anomalous energy is estimated reliably as there is a good agreement between our model and lattice data for the average gluon momentum fraction. The differential and total cross sections for $J/\psi$ photoproduction were computed and compared with the latest JLab data from GlueX and $J/\psi$‑007 experiments, showing good overall agreement. Finally, we presented predictions for $\Upsilon$ photoproduction, as the heavy-quark limit is more suitable to employ here. These predictions can be tested at future EIC and EicC.

### v2

@@ III Gluon GPDS in the Spectator Model
In this section, we present an analytical calculation of the gluon GPDs $H^{g}(x,\xi,t)$ and $E^{g}(x,\xi,t)$ at nonzero skewness $\xi$ in the proton using a spectator model [68]. A variant of this model, incorporating a spectral function for the spectator mass, has previously been used to compute twist-2 gluon TMD distributions [69, 70]. For the following discussion, it is convenient to introduce the polarization states of the proton in an arbitrary direction:
@@
The gluon distributions in $x$ satisfy the Bose symmetry [46, 66], leading to $H(x,\xi,t)=H(-x,\xi,t)$ and $E(x,\xi,t)=E(-x,\xi,t)$. As shown in Eq. (10), the integrals of gluon GPDs over $x$ are restricted to $x\geq 0$, only the two poles at $k_{2}^{-}$ and $k_{3}^{-}$ in Eqs. (28-29) need to be considered,
@@
After selecting the appropriate helicity combinations and carrying out the $k^{-}$ integrals in Eq. (26), the twist-2 gluon GPDs can be constructed as
@@ IV.1 Parameter determination and derived quantities
The present spectator model adopted in the calculation is similar to that in Ref. [69] except we refrain to include a spectral function for the spectator mass. The model contains four free parameters: the normalization constants $\kappa_{1,2}$; the cut‑off $\Lambda_{X}$ which regulates the large $\bm{k}_{\perp}$-behavior; and the spectator mass $M_{X}$ which is taken larger than the proton mass ($M_{X}>M$). To describe the $x$-dependence of the distributions reliably, we determine these parameters by simultaneously fitting the unpolarized PDF (Eq. 39) and the helicity PDF (Eq. 40) to the next-to-leading-order (NLO) NNPDF3.1 parametrization of $f_{1}^{g}(x)$ [80] and the NLO NNPDFpol1.1 parametrization of $g_{1L}^{g}(x)$ [81] at the scale $\mu_{0}=2\;\text{GeV}$. The fitted results and their uncertainties are listed in Table 1.
@@
Figure 6: Differential cross sections for $\gamma p\rightarrow J/\psi\,p$ as functions of t-t at different average photon energies $8.2\;\text{GeV}\leq E_{\gamma}\leq 11.8\;\text{GeV}$ compared with data from GlueX19 [10] (triangles), GlueX23 [11] (circles), and $J/\psi$-007 [12] (squares) . The bands depict the parameter uncertainties.
@@
We first analyze all available data for near-threshold $J/\psi$ photoproduction from the GlueX [10, 11] and $J/\psi$-007 experiments [12]. These data include differential cross sections in 15 photon-energy bins and total cross sections, with the GlueX23 measurements on differential cross section covering the full near-threshold kinematic region. They have been analyzed in recent theoretical studies on gluonic properties of the proton [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. For the parameters related to the $J/\psi$ production [18], we adopt $\alpha_{\text{EM}}=1/137$, $\alpha_{S}=0.3$, $e_{c}=2/3$, $M=0.938\;\text{GeV}$, $M_{J/\psi}=3.097\;\text{GeV}$, and the $J/\psi$ wave function at the origin [84, 85]
@@
In Fig. 6 we plot the model results (solid curves) for the differential cross sections as functions of t-t at different $\langle E_{\gamma}\rangle$ compared with the experimental data measured at different t-t bins. The triangle, circle, and square data points corresponds to the data from GlueX19 [10] , GlueX23 [11] , and $J/\psi$-007 [12] , respectively. The blue bands depict the uncertainties from the model parameters. Overall, the model describes the data reasonably well, with slight deviations at $\langle E_{\gamma}\rangle=9.86$ and $10.82\;\text{GeV}$. The differential cross sections exhibit the characteristic diffractive peak at forward $t$ and an exponential fall-off at larger momentum transfer. The GlueX23 data (circles) span the entire physical $t$ range; at the lowest average energy $\langle E_{\gamma}\rangle=8.93\;\text{GeV}$ an enhancement appears at the largest $|t|$, possibly signaling contributions beyond the pure $t$-channel mechanism (e.g., $u$- or $s$-channel exchanges) [23, 86]. While our model does not reproduce this enhancement at $8.93\;\text{GeV}$, similar features emerge at $\langle E_{\gamma}\rangle=9.86$ and $10.82\;\text{GeV}$.
@@
In Fig. 8 we show the predicted differential cross sections near threshold ($\langle E_{\gamma}\rangle=57.2\;\text{GeV}$, left) and farther above threshold ($\langle E_{\gamma}\rangle=70\;\text{GeV}$, right). The total cross section is displayed in Fig. 9. The larger bottom-quark mass makes the heavy-quark limit more reliable for $\Upsilon$ production, reducing theoretical uncertainties from higher-order corrections. However, the production rate is strongly suppressed by the heavy quark mass, leading to cross sections much smaller than that for $J/\psi$. Verification of these results awaits more precise future measurements.
@@ V Summary
In this work we studied the near-threshold exclusive photoproduction of heavy quarkonia within the GPD formalism, employing a spectator model for the proton. The amplitudes of these processes are dominated by the leading moments of gluon GPDs, offering a probe of gluon distributions at large skewness $\xi$. The model parameters were determined by simultaneously fitting $f_{1}^{g}(x)$ from the NNPDF3.1 parameterization and $g_{1L}^{g}(x)$ from the NNPDFpol1.1 parameterization at $\mu_{0}=2\;\text{GeV}$. The gluon GPDs $H^{g}(x,\xi,t)$ and $E^{g}(x,\xi,t)$ at nonzero $\xi$ were calculated analytically at tree level via two-gluon exchange, separately in the regions $0\leq x\leq\xi$ and $\xi<x\leq 1$.
Several gluon-related quantities were extracted and compared with recent lattice results. The gluonic contribution to the quantum anomalous energy is estimated reliably as there is a good agreement between our model and lattice data for the average gluon momentum fraction. The differential and total cross sections for $J/\psi$ photoproduction were computed and compared with the latest JLab data from GlueX and $J/\psi$-007 experiments, showing good overall agreement. Finally, we presented predictions for $\Upsilon$ photoproduction, as the heavy-quark limit is more suitable to employ here. These predictions can be tested at future EIC and EicC.